### v5

# Bounding User Contributions for User-Level Differentially Private Mean Estimation

Abstract
We revisit the problem of releasing the sample mean of bounded samples in a dataset, privately, under user-level $\varepsilon$-differential privacy (DP). We aim to derive the optimal method of preprocessing data samples, within a canonical class of processing strategies, in terms of the estimation error. Typical error analyses of such bounding (or clipping) strategies in the literature assume that the data samples are independent and identically distributed (i.i.d.), and sometimes also that all users contribute the same number of samples (data homogeneity)—assumptions that do not accurately model real-world data distributions. Our main result in this work is a precise characterization of the preprocessing strategy that gives rise to the smallest worst-case error over all datasets – a distribution-independent error metric – while allowing for data heterogeneity. We also show via experimental studies that even for i.i.d. real-valued samples, our clipping strategy performs much better, in terms of average-case error, than the widely used bounding strategy of Amin et al. (2019).
I Introduction
In this article, we work on the fundamental problem of processing bounded, potentially vector-valued samples in a dataset, for the release of a private estimate of the sample mean. In particular, we work within the framework of “user-level” differential privacy [15, 6], which is a generalization of the now widely adopted framework of differential privacy (DP) [9, 19] for the design and analysis of privacy-preserving algorithms. Loosely, user-level DP guarantees the privacy of a “user”, who could contribute more than one sample, by ensuring the statistical indistinguishability of outputs of the algorithm to changes in the user’s samples. User-level DP has practical relevance for inference tasks on real-world datasets, such as traffic datasets, datasets of user expenditures, and time series data, where different users contribute potentially different numbers of samples (data heterogeneity) [20, 17], and more recently, in federated learning (FL) applications (see, e.g., [14, Sec. 4]).
There are two key requirements of such user-level DP mechanisms for mean estimation. Firstly, the mechanisms must be designed to work with heterogeneous data. Secondly, one would like reliable reconstruction of the true sample mean, even when the data samples are non-i.i.d. (independent and identically distributed). Our focus is hence to characterize an error metric, which is independent of the underlying data distribution and can be explicitly computed and optimized, for heterogeneous data.
In this article, we confine our attention to user-level (pure) $\varepsilon$-DP algorithms for mean estimation. A key subroutine in such mechanisms [15, 7, 16, 2] is the preprocessing of the data samples for the release of an estimate of the sample mean, which requires the addition of less noise for privacy, as against releasing a noised version of the true sample mean. Such a preprocessing procedure (or “bounding” or “clipping” strategy [2]) either drops selected samples, or projects the samples to a “high-probability subset”. While it is usually easy to establish that such mechanisms are differentially private, an analysis of their “utility”, or the error in estimation of the true statistic, often relies on distributional assumptions about the dataset.
In this work, following [13, 3, 18], we define and explicitly compute the worst-case error, over all datasets, of general preprocessing (or bounding) strategies. This error metric is natural in settings with arbitrarily correlated data, where a user potentially ascribes his/her error tolerance to the worst dataset that the statistic is computed on. Furthermore, this error metric is distribution-independent and, as we show, is computable under data heterogeneity too. We then explicitly identify the bounding strategy that results in the smallest worst-case error. Interestingly, we also observe from experimental studies that for scalar samples, our clipping strategy also gives rise to much smaller errors on average compared to that in [2], for selected dataset sizes, when the samples are drawn i.i.d. according to common distributions.
II Notation and Preliminaries
II-A Notation
The notation $\mathbb{N}$ denotes the set of positive natural numbers. For $n\in\mathbb{N}$, the notation $[n]$ denotes the set $\{1,2,\ldots,n\}$. Further, given reals $a,b$ with $a\leq b$, we define $\Pi_{[a,b]}(x):=\min\{\max\{x,a\},b\}$, for $x\in\mathbb{R}$. For a vector $\mathbf{x}\in\mathbb{R}^{d}$, we define its $\ell_{p}$-norm $\lVert\mathbf{x}\rVert_{p}$, for an integer $p\in\mathbb{N}$ to be $(\sum_{i=1}^{d}|x_{i}|^{p})^{1/p}$, with $\lVert\mathbf{x}\rVert_{\infty}:=\max_{1\leq i\leq d}|x_{i}|$. For a given set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and a vector $\mathbf{a}$, we define, with some abuse of notation, $\mathcal{X}(\mathbf{a})$ to be an $\ell_{1}$-projection $\arg\min_{\mathbf{b}\in\mathcal{X}}\lVert\mathbf{a}-\mathbf{b}\rVert_{1}$. Given an integer $d\geq 1$ and a real $M\geq 0$, we define $\Delta_{M}$ to be the $d$-simplex, i.e., $\Delta_{M}:=\{\mathbf{a}\in\mathbb{R}^{d}:\ \sum_{i\leq d}a_{i}\leq M,\ a_{i}\geq 0,\ \text{for all $i\in[d]$}\}$, and the set $\delta_{M}:=\{\mathbf{a}\in\mathbb{R}^{d}:\ \sum_{i\leq d}a_{i}=M,\ a_{i}\geq 0,\ \text{for all $i\in[d]$}\}$. Further, for $0\leq\alpha\leq\beta$, we define, with some abuse of terminology, the “annulus” $\mathsf{A}_{\alpha,\beta}$ as $\mathsf{A}_{\alpha,\beta}:=(\Delta_{\beta}\setminus\Delta_{\alpha})\cup\delta_{\alpha}$.
We use the notation $\text{Lap}(b)$ to refer to the zero-mean Laplace distribution with standard deviation $\sqrt{2}b$, the notation $\text{Unif}((0,a])$ to denote the uniform distribution on the interval $(0,a]$, and the notation $\mathcal{N}(\mu,v)$ to denote the Gaussian distribution with mean $\mu$ and variance $v$.
II-B Problem Formulation
Let $L$ be the number of users present in the dataset. For every user $\ell\in[L]$, let the number of records contributed by the user be denoted by $m_{\ell}$, and set $m^{\star}:=\max_{\ell\in[L]}m_{\ell}$ and $m_{\star}:=\min_{\ell\in[L]}m_{\ell}$. We assume that $L$ and the collection $\{m_{\ell}:\ell\in[L]\}$ are known. Now, for a given user $\ell\in[L]$, let $\left\{\mathbf{x}_{j}^{(\ell)}:j\in[m_{\ell}]\right\}$ denote the collection of (potentially arbitrary) bounded samples contributed by the user, where each $\mathbf{x}_{j}^{(\ell)}\in\mathbb{R}^{d}$, for some dimension $d\geq 1$. We assume, as is common for most applications [1, 13], that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}$, for all $\ell\in[L]$, $j\in[m_{\ell}]$, where $U>0$ is known. Call the dataset as $\mathcal{D}=\left\{\left(\ell,\mathbf{x}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$.
We are interested in releasing the sample mean
$f=f(\mathcal{D}):=\frac{1}{\sum_{\ell^{\prime}=1}^{L}m_{\ell^{\prime}}}\cdot\sum_{\ell=1}^{L}\sum_{j=1}^{m_{\ell}}\mathbf{x}_{j}^{(\ell)}.$
II-C Differential Privacy
[…]
Definition II.1.
For a fixed $\varepsilon>0$, a mechanism $M:\mathsf{D}\to\mathbb{R}^{d}$ is user-level $\varepsilon$-DP if for every pair of user-level neighbours $\mathcal{D}_{1},\mathcal{D}_{2}$ and for every measurable subset $Y\subseteq\mathbb{R}^{d}$, we have that $\Pr[M(\mathcal{D}_{1})\in Y]\leq e^{\varepsilon}\Pr[M(\mathcal{D}_{2})\in Y].$
Definition II.2.
Given a function $g:\mathsf{D}\to\mathbb{R}^{d}$, we define its user-level sensitivity $\Delta_{g}$ as $\Delta_{g}:=\max_{\mathcal{D}_{1},\mathcal{D}_{2}\ \text{u-l nbrs.}}\lVert g(\mathcal{D}_{1})-g(\mathcal{D}_{2})\rVert_{1},$ where the maximization is over datasets that are user-level neighbours.
We use the terms “sensitivity” and “user-level sensitivity” interchangeably. The next result is well-known [8, Prop. 1].
Theorem II.1.
For any $g:\mathsf{D}\to\mathbb{R}^{d}$, the mechanism $M^{\text{Lap}}_{g}:\mathsf{D}\to\mathbb{R}$ defined by $M^{\text{Lap}}_{g}(\mathcal{D}_{1})=g(\mathcal{D}_{1})+\mathbf{Z},$ where $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$ is such that $Z_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Lap}(\Delta_{g}/\varepsilon)$, $i\in[d]$, and is independent of $\mathcal{D}$, is user-level $\varepsilon$-DP.
II-D The Worst-Case Error Metric
All through, in this paper, we shall work with user-level $\varepsilon$-DP mechanisms that add a suitable amount of Laplace noise that is tailored to the sensitivity of the function used as an estimator of the sample mean $f$. Consider a mechanism $M$ for the user-level $\varepsilon$-DP release of the statistic $f$. The canonical structure of $M$ (see [11, Footnote 1], [12]) is: $M(\mathcal{D})=\overline{f}(\mathcal{D})+{\mathbf{Z}},$ for some estimator $\overline{f}$ of $f$, with user-level sensitivity $\Delta_{\overline{f}}$, with $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$, with $Z_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Lap}\left(\Delta_{\overline{f}}/\varepsilon\right)$, $i\in[d]$.
For the mechanism $M$, we define its worst-case estimation error as
$E_{M}=E_{\overline{f}}^{(d)}:=\max_{\mathcal{D}\in\mathsf{D}}\big\lVert f(\mathcal{D})-\overline{f}(\mathcal{D})\big\rVert_{1}+\mathbb{E}[\lVert{\mathbf{Z}}\rVert_{1}].$ (1)
Clearly, the expression above is an upper bound on the $\ell_{1}$-error $E^{(1)}_{M}:=\max_{\mathcal{D}\in\mathsf{D}}\mathbb{E}[|f(\mathcal{D})-M(\mathcal{D})|]$, via the triangle inequality. When the dimension $d$ is clear from the context, we simply denote $E_{\overline{f}}^{(d)}$ as $E_{\overline{f}}$. The distribution-independent expression in (1) conveniently captures the errors due to bias (the first term) and due to noise addition for privacy (the second term); a similar such error measure that separates the bias and noise errors was employed in [2].
III Worst-Case Errors of Bounding/Clipping Strategies
III-A On Clipping Strategies
We work with estimators $\overline{f}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ of $f$ obtained by bounding user contributions as follows: for each $\ell\in[L]$ and $j\in[m_{\ell}]$, we let $\overline{\mathbf{x}}_{j}^{(\ell)}:={\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}}\left(\mathbf{x}_{j}^{(\ell)}\right)$, for reals $0\leq a_{j}^{(\ell)}\leq b_{j}^{(\ell)}\leq U$. In words, $\overline{\mathbf{x}}_{j}^{(\ell)}$ is a projection of $\mathbf{x}_{j}^{(\ell)}$ onto the set $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$ (see Figure 1), which, intuitively, reduces the range of values that $\mathbf{x}_{j}^{(\ell)}$ can take, and hence its sensitivity too. We then set
$\overline{f}=\overline{f}(\mathcal{D}):=\frac{1}{\sum_{\ell^{\prime}=1}^{L}m_{\ell^{\prime}}}\cdot\sum_{\ell=1}^{L}\sum_{j=1}^{m_{\ell}}\mathbf{\overline{x}}_{j}^{(\ell)}.$
[…]
Figure 1: The annulus $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, for $d=2$, shown in blue. Here, the points $q_{1}$, $q_{2}$ equal $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(p_{1})$ and $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(p_{2})$, respectively.
Note that the class $\mathsf{B}$ of estimators $\overline{f}$ as above captures those estimators obtained by dropping selected samples $\mathbf{x}_{j}^{(\ell)}$ (by setting $a_{j}^{(\ell)}=b_{j}^{(\ell)}=0$ for those samples) and those obtained by projecting samples onto an $\ell_{1}$-bounded subset of $\Delta_{U}$, in addition to strategies that perform a combination of dropping and projection. This class of estimators hence includes several common estimators of the sample mean used in works such as [15, 3, 18].
III-B On Worst-Case Errors of Clipping Strategies
Consider the quantity $E_{\overline{f}}^{(d)}$, for $\overline{f}\in\mathsf{B}$. The following proposition then holds:
Proposition III.1.
We have that
$E_{\overline{f}}^{(1)}$ $=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right)\ +$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}$
and for any $d\geq 2$,
$E_{\overline{f}}^{(d)}$ $=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right)\ +$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{2d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}b_{j}^{(\ell)}}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}.$
The proof of Proposition III.1 proceeds with help from a few lemmas. For $E_{\overline{f}}$ as in (1), we define $\beta^{(d)}(\overline{f})=\beta(\overline{f})$ to be the bias, i.e., $\beta(\overline{f}):=\max_{\mathcal{D}\in\mathsf{D}}\big\lVert f(\mathcal{D})-\overline{f}(\mathcal{D})\big\rVert$ and $\eta^{(d)}(\overline{f})=\eta(\overline{f})$ to be the error due to noise addition, i.e., $\eta(\overline{f}):=\mathbb{E}[\lVert\mathbf{\mathbf{Z}}\rVert_{1}]$. First, we aim to characterize $\beta(\overline{f})$. To this end, we first state a necessary condition for a vector $\mathbf{y}$ to be an $\ell_{1}$-projection of a vector $\mathbf{a}\in\Delta_{U}$ onto $\Delta_{\alpha}$, for $\alpha\leq U$.
Lemma III.1.
Given $\mathbf{a}\in\Delta_{U}$ and $\alpha\leq U$, we have $\Delta_{\alpha}(\mathbf{a})=\mathbf{a}$, if $\mathbf{a}\in\Delta_{\alpha}$. Else, any $\ell_{1}$-projection $\mathbf{y}=\Delta_{\alpha}(\mathbf{a})$ must satisfy $\lVert\mathbf{y}\rVert_{1}=\alpha$, with $y_{i}\leq a_{i}$, for all $i\in[d]$.
Proof.
The first statement in the lemma is clear. Now, for $\mathbf{a}\notin\Delta_{\alpha}$, suppose that $\mathbf{y}=\Delta_{\alpha}(\mathbf{a})$ is such that $\lVert\mathbf{y}\rVert_{1}<\alpha$. It can then be seen that by setting $\mathbf{y}^{\prime}:=\lambda\mathbf{y}+(1-\lambda)\mathbf{a}$, for some $\lambda\in(0,1)$ such that $\lVert\mathbf{y}^{\prime}\rVert_{1}=\alpha$, we will obtain that $\lVert\mathbf{y}^{\prime}-\mathbf{a}\rVert_{1}<\lVert\mathbf{y}-\mathbf{a}\rVert_{1}$, which is a contradiction. Likewise, suppose that $y_{i}>a_{i}$, for some $i\in[d]$. Now, consider any coordinate $j\in[d]$ such that $y_{j}\leq a_{j}$ (such a coordinate must exist, since $\lVert\mathbf{y}\rVert_{1}=\alpha$); by letting $m:=\min\{|y_{i}-a_{i}|,|y_{j}-a_{j}|\}$ and setting $y_{i}\leftarrow y_{i}-m$ and $y_{j}\leftarrow y_{j}+m$, we see that we strictly decrease $\lVert\mathbf{y}-\mathbf{a}\rVert_{1}$, which is a contradiction. ∎
In the appendix, we explicitly characterize the vectors $\mathbf{y}\in\Delta_{\alpha}$ that are the $\ell_{1}$-projections of $\mathbf{a}\in\Delta_{U}\setminus\Delta_{\alpha}$, for $\alpha\leq U$; indeed, we have that the condition stated in Lemma III.1 is both necessary and sufficient. The next lemma exactly characterizes the bias $\beta^{(d)}(\overline{f})$, for any $d\geq 1$.
Lemma III.2.
We have that for any $d\geq 1$,
$\beta^{(d)}(\overline{f})=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right).$
Proof.
For a sample $\mathbf{x}_{j}^{(\ell)}\in\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, it is clear that $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})=\mathbf{x}_{j}^{(\ell)}$. Now, suppose that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}\setminus\Delta_{b_{j}^{(\ell)}}$. Following Lemma III.1, if $\mathbf{y}=\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{a})$, we must have
$\lVert\mathbf{y}-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=\sum_{i\leq d}((x_{j}^{(\ell)})_{i}-y_{i})=\lVert\mathbf{x}_{j}^{(\ell)}\rVert_{1}-b_{j}^{(\ell)}.$
Hence, the worst-case clipping error for sample $\mathbf{x}_{j}^{(\ell)}$ is
$\max\ \lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=U-b_{j}^{(\ell)}.$
By symmetric arguments, one can show that if $\mathbf{x}_{j}^{(\ell)}\in\Delta_{a_{j}^{(\ell)}}\setminus\delta_{a_{j}^{(\ell)}}$, we must have
$\max\ \lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=a_{j}^{(\ell)}.$
Thus, overall, we obtain that
$\max_{\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}}\lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}.$ (2)
Now, for any dataset $\mathcal{D}$, recall that
$\lVert f(\mathcal{D})-\overline{f}(\mathcal{D})\rVert_{1}$
$=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\left\lVert\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(\mathbf{x}_{j}^{(\ell)}-\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})\right)\right\rVert_{1}.$ (3)
Putting together (2) and (3) concludes the proof. ∎
The calculation of $\eta(\overline{f})$ is quite similar to the proof above, and is captured in Lemma III.3 below.
Lemma III.3.
We have that
$\eta^{(1)}(\overline{f})=\frac{d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}$
and for any $d\geq 2$,
$\eta^{(d)}(\overline{f})=\frac{2d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}b_{j}^{(\ell)}}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}.$
Proof.
Consider first the case when $d=1$. Via arguments entirely analogous to that in the proof of Lemma III.2, the user-level sensitivity of $\overline{f}$ is
$\Delta_{\overline{f}}=\frac{\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}},$
since in the worst-case, all samples of a user $\ell$ are changed each from $(a_{j}^{(\ell)},0,\ldots,0)$ to $(b_{j}^{(\ell)},0,\ldots,0)$. Next, for any $d\geq 2$, we have via similar arguments that
$\Delta_{\overline{f}}=\frac{2\max_{\ell\leq L}\sum_{j\leq m_{\ell}}b_{j}^{(\ell)}}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}},$
since in the worst-case, all samples of a user $\ell$ are changed each from $(b_{j}^{(\ell)},0,0,\ldots,0)$ to $(0,b_{j}^{(\ell)},0,\ldots,0)$. Using $\mathbb{E}[\lVert\mathbf{Z}\rVert_{1}]=d\mathbb{E}[|Z_{1}|]=d\Delta_{\overline{f}}/\varepsilon$ gives us the required result. ∎
The proof of Proposition III.1 then follows directly by putting together Lemmas III.2 and III.3.
Given the characterization of the worst-case error $E_{\overline{f}}$ as above, we proceed with identifying an estimator $f^{\star}\in\mathsf{B}$ (equivalently, a bounding strategy) that minimizes $E_{\overline{f}}$, over all $\overline{f}\in\mathsf{B}$. Let $T_{\varepsilon}$ denote the $\left\lceil\left(\frac{2d}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest value in the collection $\{Um_{1},Um_{2},\ldots,Um_{L}\}$; if $\varepsilon<2d/L$, we set $T_{\varepsilon}=0$. Our main result is encapsulated in the following theorem:
Theorem III.1.
We have that $f^{\star}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizes $E_{\overline{f}}$, where $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{Um_{\ell}+T_{\varepsilon}}{2m_{\ell}},U\right\},$ if $d=1$, and $a_{j}^{(\ell)}=0\ \text{ and }b_{j}^{(\ell)}=\min\left\{\frac{T_{\varepsilon}}{m_{\ell}},U\right\}$, for any $d\geq 2$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
Some remarks are in order. First, note that the optimal bounding strategy $f^{\star}$ clips sample values based only on the number of contributions of each user $\ell\in[L]$. Furthermore, note that the interval of projection is determined by $T_{\varepsilon}$, which is quite similar in structure to the optimal clipping threshold $T$ in [2] for item-level DP, which is the (privately estimated) $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample value.
The proof of Theorem III.1 proceeds with help from the following lemma.
Lemma III.4.
For $d=1$, there exists an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$, which obeys $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$. Furthermore, for $d\geq 2$, an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$ must set $a_{j}^{(\ell)}=0$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
Proof.
Consider any optimal estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$, and suppose that $a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})}>U$, for some $\tilde{\ell}\in[L]$ and $j\in[m_{\tilde{\ell}}$]. The proof when we suppose that $a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})}<U$, for some $\tilde{\ell}\in[L]$ and $j\in[m_{\tilde{\ell}}$], is similar, and is omitted. Let $\delta=(a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})})-U$. By setting $b_{j}^{(\tilde{\ell})}\leftarrow b_{j}^{(\tilde{\ell})}-\delta$, we observe that $\beta^{(d)}(\overline{f})$ remains unchanged, while $\eta^{(d)}(\overline{f})$ either remains unchanged or strictly decreases by $\delta>0$. Moreover, for $d\geq 2$, note from Proposition III.1 that increasing any $a_{j}^{(\ell)}$ from 0 to a positive value can only strictly increase $E_{\overline{f}}$. ∎
Hence, to obtain the explicit structure of an estimator $\overline{f}$ that minimizes $E_{\overline{f}}$, it suffices to focus estimators $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ with $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$, $j\in[m_{\ell}]$. Then, for $d=1$, there exists an estimator $f^{\star}$ minimizing $E^{(1)}_{\overline{f}}$ that satisfies
$E^{(1)}_{f^{\star}}=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{1}{\varepsilon}\cdot\max_{\ell\leq L}\left(Um_{\ell}-2S^{(\ell)}\right)\right),$ (4)
where for $\ell\in[L]$, we let $S^{(\ell)}:=\sum_{j\leq m_{\ell}}a_{j}^{(\ell)}$, for any given estimator $\overline{f}$. Furthermore, for any $d\geq 2$, there exists an estimator $f^{\star}$ minimizing $E^{(d)}_{\overline{f}}$ that satisfies
$E^{(d)}_{f^{\star}}=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\overline{S}^{(\ell)}+\frac{2d}{\varepsilon}\cdot\max_{\ell\leq L}\left(Um_{\ell}-\overline{S}^{(\ell)}\right)\right),$ (5)
where for $\ell\in[L]$, we let $\overline{S}^{(\ell)}:=\sum_{j\leq m_{\ell}}\left(U-b_{j}^{(\ell)}\right)$, for any given estimator $\overline{f}$. We now prove Theorem III.1.
Proof of Thm. III.1.
We prove the theorem in detail for the case when $d=1$; the setting of $d\geq 2$ is quite similar and the proof is hence is omitted. We begin with the expression in (4); note that our task now is simply to identify the optimal parameters $\{S^{(\ell)}:\ \ell\in[L]\}$ of $f^{\star}$ in (4). Once these parameters are derived, we simply set $a_{j}^{(\ell)}:=S^{(\ell)}/m_{\ell}=U-b_{j}^{(\ell)}$ (via Lemma III.4), for each $\ell\in[L]$ and $j\in[m_{\ell}]$. In the expression in (4), let us set $\tau:=\max_{\ell\leq L}\left(Um_{\ell}-2S^{(\ell)}\right)$. We hence need to solve the following constrained optimization problem:
${\text{minimize}}\quad h(\{S^{(\ell)}\}):=\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{\tau}{\varepsilon}\right)$
$\text{subj. to:}\ \ Um_{\ell}-2S^{(\ell)}\leq\tau,\ S^{(\ell)}\geq 0,\ \forall\ \ell\in[L],\ \tau\geq 0.$ (6)
The optimization problem in (6) is a linear programming problem. By standard arguments via the necessity of the KKT conditions [4, Sec. 5.5.3], there must exist reals $\lambda_{\tau}\geq 0$ and $\lambda_{\ell},\ \mu_{\ell}\geq 0$, for each $\ell\in[L]$ (or Lagrange multipliers), such that the function
$\mathcal{L}(\{S^{(\ell)}\},\$$\lambda_{\tau},\{\lambda_{\ell},\mu_{\ell}\}):=\sum_{\ell\leq L}S^{(\ell)}+\frac{S}{\varepsilon}-\lambda_{\tau}\tau\ +$
$\sum_{\ell}\lambda_{\ell}\cdot\left(Um_{\ell}-2S^{(\ell)}-S\right)-\sum_{\ell}\mu_{\ell}S^{(\ell)}$
obeys the following properties.
  • Stationarity: We have that $\frac{\partial\mathcal{L}}{\partial\tau}=0$, or $\lambda_{\tau}+\sum_{\ell}\lambda_{\ell}=\frac{d}{\varepsilon},$ and that $\frac{\partial\mathcal{L}}{\partial S^{(\ell)}}=0$, for each $\ell\in[L]$, or $\lambda_{\ell}=\frac{1-\mu_{\ell}}{2}.$
  • Complementary slackness: We have that
    $\lambda_{\tau}\tau=0,\ \lambda_{\ell}\cdot\left(Um_{\ell}-2S^{(\ell)}-\tau\right)=0,\ \text{and}\ \mu_{\ell}S^{(\ell)}=0,$
    for all $\ell\in[L]$.
We claim that the assignment $\tau^{\star}=T_{\varepsilon}$ and $S^{(\ell),\star}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2},0\right\}$ satisfies the conditions above, for an appropriate choice of $\lambda_{\tau},\{\lambda_{\ell},\mu_{\ell}\}$ values. Indeed, for $\ell\in[L]$, if $S^{(\ell),\star}=0$, we set $\lambda_{\ell}=0$ and $\mu_{\ell}=1$; else, we set $\lambda_{\ell}=\frac{1}{2}$ and $\mu_{\ell}=0$.
Observe that from Theorem III.1, the optimal worst-case error when $d=1$ is
$E^{\text{OPT, $(1)$}}(\varepsilon)=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2},0\right\}+\frac{T_{\varepsilon}}{\varepsilon}\right).$ (7)
and the the optimal worst-case error for any $d\geq 2$ is such that
$E^{\text{OPT, $(d)$}}(\varepsilon)$
$=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{{Um_{\ell}-T_{\varepsilon}},0\right\}+\frac{2dT_{\varepsilon}}{\varepsilon}\right).$ (8)
We end this section with a couple of remarks. First, observe from (7) and (8) that, in the limit as $\varepsilon\to 0$, the optimal worst-case error equals the worst-case bias error, which in turn equals $U/2$, for $d=1$, and $U$, when $d\geq 2$.
Next, consider the special case of Theorem III.1 when $d=1$, with the interpretation that each sample $x_{j}^{(\ell)}\in[0,U]$. Here, the annulus $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$ is simply the interval $a_{j}^{(\ell)},b_{j}^{(\ell)}]\subseteq[0,U]$. This implies that for the case when the samples $\mathbf{x}_{j}^{(\ell)}$ are allowed to take values in the cube $0,U]^{d}$, as against in the $\ell_{1}$-ball $\lVert\mathbf{x}_{j}^{(\ell)}\rVert_{1}\leq U$, one can perform the bounding procedure, using the same interval $a_{j}^{(\ell)},b_{j}^{(\ell)}]$ identified via Theorem III.1 for the $d=1$ setting, for each dimension, independently.
IV Numerical Experiments
In this section, we compare, via numerical experiments, the performance of our clipping strategy for the case when $d=1$, with the widely used strategy in [2, Sec. 3], which we call the “AKMV” mechanism [^1] The subscript “AKMV” stands for the initials of the last names of the authors of [2].. We straightforwardly adapt the mechanism in [2] from the item-level setting to the user-level setting, by considering the contribution of a user to be the sum $\sum_{j\leq m_{\ell}}x_{j}^{(\ell)}$ of the samples it contributes in the user-level setting.
Importantly, the AKMV mechanism uses a privacy budget of $\varepsilon/2$ to first privately estimate the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample, and then uses this value as the clipping threshold. Via entirely heuristic analysis, the authors of [2] argue the “optimality” of AKMV mechanism. However, importantly, their analysis ignores the error in the private estimation of the the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample, rendering such a claim of optimality incorrect. As we shall see, the clipping strategy in [2, Sec. 3] performs quite poorly in comparison to our worst-case-error-optimal strategy, in the average case too, when the data samples are drawn i.i.d. from natural distributions.
IV-A Experimental Setup
A natural application of user-level DP mechanisms is to spatio-temporal datasets; similar to previous work [13, 3], we let $U=65$, in line with the largest speed of buses in km/hr, in Indian cities. We consider two collections of numbers of user contributions $\{m_{\ell}:\ \ell\in[L]\}$.
  1. 1.
    Geometric collection: Here, we fix an integer $M$ and consider $L=2^{M+1}-1$ users; for each $i\in\{0,1,\ldots,M\}$, we create $2^{i}$ users each contributing $2^{M-i}$ samples. In our experiments, we set $M=6$.
  2. 2.
    Extreme-valued collection: Here, we consider $L$ users where L1L-1 users contribute one sample each and $1$ user contributes $m^{\star}>1$ samples. In our experiments, we pick $L=101$, with $m^{\star}=10$.
We also work with the following synthetically generated datasets with i.i.d. samples:
  1. 1.
    Uniform samples: Each sample $x_{j}^{(\ell)}\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\text{Unif}((0,U])$, across $\ell\in[L]$, $j\in[m_{\ell}]$.
  2. 2.
    Projected Gaussian samples: Each sample $x_{j}^{(\ell)}$, $\ell\in[L]$, $j\in[m_{\ell}]$, is drawn i.i.d. by rejection sampling from the $\mathcal{N}(U/2,U/4)$ distribution so that the samples lie in $(0,U]$.
IV-B Performance on Synthetic Datasets
We compare the performances of the following three mechanisms on i.i.d. synthetic datasets:
  1. 1.
    the vanilla Laplace mechanism that releases $M_{\text{Lap}}(\mathcal{D})=f(\mathcal{D})+\mathbf{Z}_{1}$, where $\mathbf{Z}_{1}\sim\text{Lap}\left(\frac{\Delta_{f}}{\varepsilon}\right)$;
  2. 2.
    the “OPT-worst-case” mechanism that releases $M_{\text{OPT-wc}}(\mathcal{D})=f^{\star}(\mathcal{D})+\mathbf{Z}_{2}$, where $\mathbf{Z}_{2}\sim\text{Lap}\left(\frac{\Delta_{f^{\star}}}{\varepsilon}\right)$; and
  3. 3.
    the “AKMV” mechanism [2, Sec. 3] that releases $M_{\text{AKMV}}(\mathcal{D})=\widehat{f}_{T}(\mathcal{D})+\mathbf{Z}_{3}$, where $\mathbf{Z}_{3}\sim\text{Lap}\left(\frac{2\Delta_{\widehat{g}_{T}}}{\varepsilon}\right)$, and $\widehat{f}_{T}(\mathcal{D})$ is obtained by clipping the sum of samples of each user to lie in $[0,T]$. Here, $T$ is the estimate of the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample among $\{\sigma^{(\ell)}\}$, which is estimated with privacy budget $\varepsilon/2$.
The“ average-case” errors on i.i.d. datasets are defined as
$\overline{E}_{\text{Lap}}=\mathbb{E}[|\mathbf{Z}_{1}|],\ \overline{E}_{\text{OPT-wc}}=\mathbb{E}[|M_{\text{OPT-wc}}(\mathcal{D})-f(\mathcal{D})|],$
and
$\overline{E}_{\text{AKMV}}=\mathbb{E}[|M_{\text{AKMV}}(\mathcal{D})-f(\mathcal{D})|],$
where the expectations are over the randomness in the data samples and in the DP mechanism employed. These errors are then estimated via Monte-Carlo averaging over $10^{4}$ iterations. As a pre-processing step, we replace each of the samples $\{x_{j}^{(\ell)}:\ j\in[m_{\ell}]\}$ of every user $\ell\in[L]$ by the sample average $\frac{1}{m_{\ell}}\cdot\sum_{j\leq m_{\ell}}x_{j}^{(\ell)}$, so as to allow for improved performance post clipping.
Figure 2 and 3 (with the error axis on a log-scale), respectively, show plots of the (estimates of) the average-case errors for the three mechanisms above for the geometric collection of $\{m_{\ell}\}$ values and uniform samples, and for the extreme-valued collection with projected Gaussian samples. Interestingly, the average-case performance of the AKMV mechanism is quite similar to, and a little worse than, the average-case performance of the vanilla Laplace mechanism, which in turn is significantly worse than the average-case error incurred by the OPT-worst-case mechanism. A reason for the poor performance of the AKMV mechanism overall is due to the reduced privacy budget allocated to the private release of the clipped estimator. Moreover, in contrast to the OPT-worst-case mechanism, the AKMV mechanism requires the private estimation of the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$ quantile value, which incurs additional error. We expect that similar trends can be observed for i.i.d. datasets with samples drawn from other distributions of $\{m_{\ell}\}$ values and samples. We add that while the clipping error remains roughly the same for small changes in $\varepsilon$, there exist certain values of $\varepsilon$ that give rise to sharp discontinuities in the $T_{\varepsilon}$ values; such behaviour hence results in the (estimates of the) average-case errors not being monotonic in $\varepsilon$, in Figure 2.
Figure 2: Average-case errors using a geometric collection of $\{m_{\ell}\}$ values and uniform samples
Figure 3: Average-case errors using an extreme-valued collection of $\{m_{\ell}\}$ values and projected Gaussian samples
V Conclusion
In this paper, we revisited the fundamental problem of releasing the sample mean under user-level differential privacy (DP) – a problem that is well-studied in the DP literature, but typically in the presence of additional (e.g., i.i.d.) assumptions on the distribution of data samples. In our work, we make no distributional assumptions on the dataset; instead, we adopt a worst-case approach to studying the estimation error. We then explicitly characterize this error for a broad class of strategies for bounding user contributions by some combination of clipping the numbers of user contributions and/or clipping the sample values themselves. Our analysis allows us to identify the bounding strategy that is optimal with respect to our worst-case error metric. Via numerical experiments, we demonstrate that our strategy is not only optimal for the worst-case error, but also performs much better than the well-known strategy in [2] for datasets with i.i.d. samples.
An interesting line of future work will be to extend our worst-case error analysis to the user-level DP release of other statistics, and identify the optimal bounding strategies in those cases as well.
Appendix A Characterizing $\ell_{1}$-Projections Onto $\Delta_{\alpha}$
In this section, we argue that the condition in the second statement of Lemma III.1 is sufficient for the vector $\mathbf{y}$ to be an $\ell_{1}$-projection of $\mathbf{a}\in\Delta_{U}\setminus\Delta_{\alpha}$, for $\alpha\leq U$.
Following Lemma III.1, consider the collection $\mathcal{S}\subseteq\mathbb{R}^{d}$ of points defined as $\mathcal{S}:=\{\mathbf{z}\in\delta_{\alpha}:\ z_{i}\leq a_{i},\ \text{for all $i\in[d]$}\}$. Note that when $a_{i}>\alpha$, for all $i\in[d]$, we have $\mathcal{S}=\delta_{\alpha}$. Observe that $\mathcal{S}$ is a convex subset of $\mathbb{R}^{d}$; hence, by a version of Carathéodory’s Theorem (see the remark after [5, Thm. 15.3.5] and [10]), any point $\mathbf{z}$ in $\mathcal{S}$ can be written as a convex combination of finitely many, in particular, $d$ points in $\mathcal{S}$. Hence, consider any such collection $\mathcal{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{d}\}$ whose convex hull equals $\mathcal{S}$. The following claim then holds.
Proposition A.1.
For any $\mathbf{z}\in\mathcal{S}$, we have that $\lVert\mathbf{a}-\mathbf{z}\rVert_{1}$ is a constant.
Proof.
Recall that by the version of Carathéodory’s Theorem above, any point $\mathbf{z}\in\mathcal{S}$ can be written as $\mathbf{z}=\sum_{k=1}^{d}\lambda_{k}\mathbf{z}_{k}$, where $\lambda_{k}\geq 0$, with $\sum_{k\leq d}\lambda_{k}=1$. It suffices to show that $\lVert\mathbf{a}-\mathbf{z}\rVert_{1}$ is independent of $\{\lambda_{k}\}$. To see this, note that
$\lVert\mathbf{a}-\mathbf{z}\rVert_{1}$ $=\sum_{r\leq d}(a_{r}-z_{r})$
$=\sum_{r\leq d}(a_{r}-\sum_{k\leq d}\lambda_{k}z_{k,r})$
$=\sum_{r\leq d}a_{r}-\sum_{k\leq d}\lambda_{k}\cdot\sum_{r\leq d}z_{k,r}=\lVert\mathbf{a}\rVert_{1}-\alpha,$
which is independent of $\{\lambda_{k}\}$. ∎
We hence have that the condition in Lemma III.1 is sufficient for the vector $\mathbf{y}$ in its statement to be an $\ell_{1}$-projection of $\mathbf{a}$. In practice, given a vector $\mathbf{a}\in\Delta_{U}\setminus\Delta_{\alpha}$, one can use the vector $\mathbf{y}=\frac{\alpha}{\lVert\mathbf{a}\rVert_{1}}\mathbf{a}$ as an $\ell_{1}$-projection.